OpenStudy (anonymous):

Prove that the fraction 21n+4/14n+3 is irreducible for every natural number n.

OpenStudy (anonymous):

dm h to slv kro

OpenStudy (anonymous):

yeh question me hi dal to better rahega

OpenStudy (anonymous):

qstn ki value km ho jaegi

OpenStudy (anonymous):

(7*3n+4)/7*2n+3 =3n+(4/7)/2n+(3/7) since 4/7&3/7 are irrational so it is unreducible . aur kuch chahie?//

OpenStudy (anonymous):

dum hai to dekh

OpenStudy (anonymous):

kya hua :)

OpenStudy (anonymous):

let me see 1st yaar

OpenStudy (anonymous):

If you can make a linear combination of the numerator and denominator equal to 1, then the numbers are relatively prime.

OpenStudy (anonymous):

ye kya hai bhai! if u ad irrat. to rat. , u get an irrat. and if you divd irrat by irrat u get irrat till nume. and denom. are not same yeh solution nahi hai, it dsnt prove anything

OpenStudy (anonymous):

my response was to 009infinity.

OpenStudy (anonymous):

please present me wth solution bt not hints.

OpenStudy (anonymous):

U may look for it yourself, it is not too hard to see it in this case....

OpenStudy (anonymous):

irt. +rat./irt.+rat.=irt/irt which is unreducible

OpenStudy (anonymous):

my maths is very X 3 weak....

OpenStudy (anonymous):

pahle khud samajh le thik se ek quest ko ek quest ko bahut tarike se kar sakte hain

OpenStudy (anonymous):

i m talking about this quest

OpenStudy (anonymous):

You want s and t such that s*(14n +3) + t*(21n +t) = 1 If you can find such s and t, then the numerator and denominator are relatively prime and so the fraction is irreducible.

OpenStudy (anonymous):

bahut tariko mein galat bhi bahut hotehain jaise ye

OpenStudy (anonymous):

maf kar solution bata

OpenStudy (anonymous):

well i'll look for that.....thnx estudier

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

baat maafi ki nhi hai u have tried efficiently and i appriciate that bt u r xplting a basic princpl...... sln mughe bhi nhi pata

OpenStudy (anonymous):

which principle plz tell i will be thankfull

OpenStudy (anonymous):

tu web pr nhi mane ga scool me baat krenge wha aur bhi honge....

OpenStudy (anonymous):

bata de plz gyan baatna chahie

OpenStudy (anonymous):

gyan btta nhi, grahan kiya jata hai

OpenStudy (anonymous):

abe bhai bata de ab kya pair padu

OpenStudy (anonymous):

3/7 and 4/7 r nt irrational.....basic prncipl violated..... vn nobdy is wathing u....... us calculator

OpenStudy (anonymous):

ok thik hai.....:)

OpenStudy (anonymous):

yaar itni der tak type karne ke baad itna likhta hai........tu to mughse bhi gaya guzra hai ;)

OpenStudy (anonymous):

message padh le